OpenStudy (anonymous):

Need help The average cricket jumps vertically with an inital upward speed of 10 ft/s. What is the hang time of such a jump, ignoring air resistance? Use the formula h=-16t^2+10t, where h is the height of the cricket in feet and t is the time in seconds after the jump. Round your answer to the nearest tenth. A. 0.4 seconds B. 0.6 seconds C. 0.8 seconds D. 1.0 second

OpenStudy (anonymous):

am saying A but not sure or B

OpenStudy (compassionate):

Hi, What would you define as, "Hang time."

OpenStudy (anonymous):

time in the air

OpenStudy (compassionate):

So you know it travels at 10 ft/s You're given the formula h=-16t^2+10t What do you think you should do? HINT: The Quadratic Formula

OpenStudy (anonymous):

IT has to be A

OpenStudy (anonymous):

we calculate the time when h=0 i.e.,initially position and again when it comes on the ground. \[-16 t^2+10t=0,-2t \left( 8t-5 \right)=0\] either -2t=0 \[\implies t=0\] it gives initial time. or 8 t-5=0 \[t=\frac{ 5 }{ 8 }=0.625\] it gives final time. so it is in air ?